### v1

# Performance Comparison of HTTP/3 and HTTP/2: Proxy vs. Non-Proxy Environments

Abstract.
This paper provides a systematic evaluation of the performance of QUIC/HTTP3(H3) and TCP/HTTP2(H2) protocols in proxy-enhanced environments. By leveraging features such as UDP-based flow-controlled streams, integrated TLS, multiplexed connections, and connection migration, H3 promises enhanced web communication. Despite extensive research, the impact of proxy integration and connection migration remains underexplored. This study addresses this gap by evaluating these protocols across various scenarios in noisy networks and proxy setups. Our findings reveal that H3 excels under high loss and latency conditions, significantly benefiting from its connection migration and multiplexing features. H3’s connection migration remains robust, maintaining stable performance even in proxy-enhanced environments, ensuring seamless network transitions. The proxy has a more neutral impact on H3, while it significantly enhances H2 performance, especially when using BBR. Any improvements observed in H3 under a proxy are minor and do not fundamentally alter H3’s performance as they do for H2. Importantly, while H2 with the right congestion control algorithm (CCA) can achieve performance comparable to H3, H3’s performance is more robust, as it is less impacted by network conditions, proxy settings, and CCA variations.
QUIC, HTTP, Multiplexing, Connection Migration, Proxy, Web Performance
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Networks Transport protocolsccs: Networks Network performance analysisccs: Networks Network measurement
1. Introduction
The rapid evolution of network technologies, driven by the increasing demands of internet-based services, has led to significant advancements in communication protocols and network architecture. Among these advancements, the development of QUIC (Iyengar and Thomson, 2022) and HTTP/3 (Bishop, 2022) significantly improve the speed, reliability, and security of web communications. QUIC, initially developed by Google and later standardized by the IETF(Wik, 2024), serves as the foundation for HTTP/3 and has the potential to reshape the Internet and its applications and services.
QUIC introduces several innovative features that distinguish it from previous protocols. It enables faster connection establishment and supports multiplexed connections without head-of-line (HoL) blocking, a substantial limitation in TCP. Furthermore, QUIC uniquely supports connection migration, which allows ongoing sessions to continue even when the client’s network endpoint changes (e.g., when switching from Wi-Fi to cellular networks) (Iyengar and Thomson, 2022). HTTP/3 builds on QUIC by incorporating these advancements to streamline data transmission processes, enhancing web page load times and responsiveness through smarter packet processing and prioritization of requests (Bishop, 2022).
Major companies such as Google (chr, 2013), Akamai (Bishop, 2021), Meta (Joras and Chi, 2022), Microsoft (Banks, 2024), and Cloudflare (Ghedini and Lalkaka, 2019) have adopted QUIC commercially, indicating its wide adoption and significant impact on the global web infrastructure. As of April 1, 2024, 71.4% of desktop browsers and 70.8% of mobile browsers support H3 among the top 1000 websites (htt, 2024), underscoring its growing relevance.
Despite prior research exploring the performance of QUIC (Ganji and Shahzad, 2021; Kakhki et al., 2017; Marx et al., 2020; Piraux et al., 2018; Shreedhar et al., 2022; Wolsing et al., 2019; Yu and Benson, 2021; Trevisan et al., 2021; Jaeger et al., 2023; Rüth et al., 2019; Wang et al., 2018; Liu and Crowley, 2023; Nepomuceno et al., 2018; Bhat et al., 2017), a critical gap remains in understanding how these protocols perform in the presence of proxies. Proxies are ubiquitous in modern web architectures, serving essential roles in content delivery, load balancing, and security. However, the interaction between these next-generation protocols and proxies is still underexplored, particularly in terms of their impact on connection migration, performance in noisy networks, and overall user experience.
This work aims to address these gaps by providing a comprehensive performance comparison of TCP/HTTP2 (H2) and QUIC/HTTP3 (H3) in both proxy and non-proxy environments, focusing on how these protocols behave under real-world impairments such as packet loss and network delay. Specifically, we measure key performance metrics, including total request time and visual metrics like Speed Index (spe, 2024), to characterize the performance of these protocols across various use cases such as single-stream, multi-stream, connection migration, and web page loading. These measurements are critical to understanding how H2 and H3 impact web performance and user-perceived experience, particularly in complex network environments involving proxies.
This paper offers several novel contributions to the understanding of H2 and H3 performance in diverse network environments, with a particular emphasis on proxy usage and connection migration:
Comprehensive Performance Evaluation with Proxy Integration: Unlike prior work, which predominantly focuses on direct connections, this paper provides the first thorough comparative analysis of H2 and H3 performance in both proxy-enhanced and non-proxy environments. Our results show that while proxies significantly enhance H2’s performance, especially when paired with the right congestion control algorithm (e.g., BBR), their impact on H3 is more neutral due to QUIC’s inherent features. This comparison sheds light on the practical implications for protocol selection in real-world proxy-based networks.
First-of-its-Kind Analysis of Connection Migration under Proxy Conditions: To the best of our knowledge, this paper is the first to evaluate QUIC’s connection migration capabilities under proxy conditions, including scenarios involving IP address changes and network impairments. Our results demonstrate H3’s ability to maintain low request times and ensure seamless user experience during network transitions, highlighting its suitability for mobile and dynamic networking environments.
Experimentation with Congestion Control Algorithms (CCAs) for H2 and H3: Through experiments with multiple CCAs (e.g., BBR, Cubic, NewReno), this study provides insights into how different CCAs affect H2 and H3 performance. We highlight conditions where BBR notably improves H2’s performance, while demonstrating that H3’s internal congestion control mechanisms ensure consistent performance across various network conditions, making it a reliable choice for broader use cases.
All our study data and scripts are available on GitHub.
2. Related Work
Performance of QUIC Extensive research has explored QUIC’s enhancements over traditional protocols, emphasizing its ability to significantly improve web browsing experiences. Early investigations by Carlucci et al. (Carlucci et al., 2015) demonstrated QUIC’s superior performance over TCP and SPDY (Wang et al., 2014), particularly in terms of connection setup times and robustness to packet loss. Cunha et al. (Da Cunha et al., 2023) further highlighted the conditions in which QUIC excels and when it underperforms. QUIC’s evolution has led to efforts investigating interoperability across implementations (Jaeger et al., 2023; Marx et al., 2020; Seemann and Iyengar, 2020). Long-term studies by Kakhki et al. (Kakhki et al., 2017) and Piraux et al. (Piraux et al., 2018) traced QUIC’s progression over time. Additionally, Rüth et al. (Rüth et al., 2018) examined QUIC deployment and usage in its early stages, while QScanner (Zirngibl et al., 2021) analyzed early QUIC deployments.
Research has also explored QUIC’s impact on various applications. For video streaming, studies by Palmer et al. (Palmer et al., 2018), Paiva et al. (do Prado Paiva et al., 2023), Zheng et al. (Zheng et al., 2021), Shreedhar et al. (Shreedhar et al., 2022), and Xu et al. (Xu et al., 2020) have highlighted QUIC’s benefits. In web browsing, research by Das et al. (Das, 2014), Rüth et al. (Rüth et al., 2019), Wolsing et al. (Wolsing et al., 2019), and Shreedhar et al. (Shreedhar et al., 2022) have shown how QUIC can enhance user experiences. Studies on mixed workloads by Shreedhar et al. (Shreedhar et al., 2022) and Wolsing et al. (Wolsing et al., 2019) extend these findings. Additionally, research has evaluated QUIC’s performance on different platforms, including mobile environments (Ganji and Shahzad, 2021) and satellite communications (Endres et al., 2022; Kosek et al., 2022b). Siag et al. (Siag et al., 2023) also developed a predictive model for QUIC’s performance. Kempf et al.(Kempf et al., 2024) presents an extension to the QUIC Interop Runner for testing the interoperability of QUIC implementations.
Features of QUIC Research on H2 and H3, such as the studies by Marx et al. (Marx et al., 2019) and Sander et al. (Sander et al., 2022), has provided in-depth analysis of how resource multiplexing and prioritization affect web performance. These studies highlight that while H3’s use of QUIC reduces HoL blocking compared to H2, real-world gains depend heavily on effective prioritization strategies. Other research has extended QUIC to multipath QUIC (MPQUIC) (Coninck and Bonaventure, 2017; Ingenzi et al., [n. d.]), and new packet schedulers for MPQUIC (do Prado Paiva et al., 2023; Rabitsch et al., 2018; Wang et al., 2019) and upper-layer solutions (Zheng et al., 2021) have followed. MultipathTester (Coninck and Bonaventure, 2019) compares the performance of MPTCP with MPQUIC on iOS. Several studies have focused on QUIC congestion control (Engelbart and Ott, 2021; Haile et al., 2022; Mishra et al., 2022), while others examine security and privacy (Kosek et al., 2022c; Lychev et al., 2015; Nawrocki et al., 2021, 2022; Soni and Rajput, 2020). Experimental protocols inspired by QUIC, like TCPLS (Rochet et al., 2021), DCQUIC (Tan et al., 2021), and MCQUIC (Navarre et al., 2023), further show its potential.
Connection Migration in QUIC Despite the importance of connection migration in QUIC for mobile and transitioning networks, detailed analyses focused on this feature in diverse networking contexts are scarce. Google has demonstrated the benefits of connection migration in a demo video (goo, 2024), but empirical studies remain limited. Tan et al. (Tan et al., 2020) discuss how QUIC uses connection migration to uniquely identify connections, proposing a Proactive Connection Migration (PCM) mechanism to optimize network selection in heterogeneous environments.
QUIC with Proxy Exploring QUIC in proxy architectures is relatively sparse, but some studies have started to bridge this gap. Kosek et al. (Kosek et al., 2023) introduced Secure Middlebox-Assisted QUIC (SMAQ), enabling middleboxes in QUIC connections while preserving security, and demonstrated performance improvements in satellite networks. Another study by Kosek et al. (Kosek et al., 2022a) evaluated QUIC and HTTP/3 with and without performance enhancing proxies (PEPs) in satellite networks, showing significant performance gains, especially in high-latency conditions. Kühlewind et al. (Kühlewind et al., 2021) assessed the QUIC-based MASQUE proxying protocol for tunneling UDP and IP traffic, revealing benefits like local recovery in lossy links despite introducing overhead. However, these studies focus on different aspects of QUIC and proxies. Our work uniquely compares H2 and H3 in proxy-enhanced environments, with a particular emphasis on QUIC’s connection migration feature, evaluating the performance impacts of proxy and connection migration under various network conditions.
Research Gap and Study Focus This paper aims to fill these gaps by systematically examining QUIC and HTTP/3’s performance across single-stream, multi-stream, connection migration, and web page loading scenarios. By focusing on the fine-grained performance aspects of QUIC and HTTP/3, especially the under-researched area of connection migration within complex architectures, this study provides nuanced insights that could guide future optimizations and implementations of these protocols in networks containing proxies. This research not only fills a knowledge gap but provides experimental design and software for future studies to explore the scalability and robustness of QUIC and HTTP/3 in dynamic and varied network settings.
Table 1. Experiment Setup
Scenarios Clients Server Network Conditions Workloads Metrics
Single Stream Request Proxygen (H3) Curl (H2, H3) Chrome (H2, H3) Wget (H2) Quiche (H3) aioquic (H3) Nginx Bandwidth: 10 Mbps Loss: 0, 2, 4(%) Delay: 0, 25, 50(ms) Small Files: 0KB, 10KB, 100KB Medium Files: 500KB, 1MB, 5MB Large Files: 10MB Total Request Time
Multi-Stream Request Curl (H2, H3) Nginx Bandwidth: 10 Mbps Loss: 0, 2, 4, 8, 12(%) Delay: 0 ms 20 Files: 1MB each Total Request Time
Connection Migration Curl (H2) Wget (H2) Quiche (H3) Nginx Bandwidth: 10 Mbps Loss: 0, 2, 4(%) Delay: 0, 25, 50(ms) One 5MB file Total Request Time
Web page Loading Chrome (H2, H3) Nginx Bandwidth: 10 Mbps Loss: 0, 2, 4, 8, 12(%) Delay: 0, 25, 50(ms) 18 popular local host websites Speed Index
3. Repeatable Experimental Setup
We use a hardware network test bed that is remotely accessible for networking researchers and educators (Wiseman et al., 2008). The client and proxy machines run Ubuntu 20.04.6 LTS on Intel(R) Xeon(R) E5520 CPUs with 16GB RAM, while the server runs Ubuntu 22.04.4 LTS on an Intel(R) Core(TM) i3-3240 CPU with 16GB RAM. These machines are interconnected through a router via a 1 Gbps Ethernet network. The ’tc netem’ utility (tc, 2024) is used for bandwidth control, packet loss, and delay manipulation. In all experiments, the bandwidth is set to 10 Mbps, with specific loss and delay parameters detailed in each experiment’s setup (see Table 1). We use a fixed bandwidth to maintain controlled conditions, allowing us to isolate the effects of proxy environments on the performance of QUIC and HTTP/2. These bandwidth limitations, combined with varied loss and delay parameters on both uplink and downlink, represent realistic edge-case network conditions that induce unique protocol behaviors for root cause analysis. The study by Zhang et al. (Zhang et al., 2024) has already explored performance under varying bandwidth, while both our work and studies like (Yu and Benson, 2021) use fixed bandwidth with varied loss and delay to focus specifically on how network impairments affect protocol performance, without introducing the additional complexity of fluctuating bandwidth.
We use Nginx (ngi, 2024) as the server and Envoy Proxy (env, 2024) as the reverse proxy, both widely adopted in practice. Six clients are employed: Proxygen (pro, 2024), Curl (cur, 2024), Chrome (chr, 2024), Wget (wge, 2024), Quiche (qui, 2024a), and aioquic (aio, 2024), ensuring a representative sample. The selection of clients, server configurations, and versions is detailed in Appendix A.1 Table 2, informed by the QUIC workgroup’s implementation list (qui, 2024b). Our criteria included support for HTTP/3, adherence to QUIC version 1, open-source availability, and popularity based on GitHub stars. Moreover, we ensured diversity by including implementations in different programming languages, aiming to reflect a broad spectrum of real-world applications.
In each experiment, we conducted 10 runs. We use Total Time as the metric in single-stream, multi-stream, and connection migration experiments, measuring from the client’s first request packet to the server until the client closes the connection. For web page loading, we use the Speed Index (spe, 2024), a metric that measures how quickly content appears on the user’s screen.
3.1. Baseline Experiment Setup
This experiment establishes the baseline for evaluating the performance of H2 and H3 in single-stream request, multi-stream request, connection migration, and web page loading. We employ Cubic as the CCA on the client side, and New Reno on the server side for both QUIC and TCP. Cubic is the default CCA for most of the H3 clients and the default CCA on Linux. For QUIC, the New Reno is the only CCA that Nginx supports. Both H2 and H3 use TLS 1.3. Our experimental setup is detailed in Table 1, with the topology illustrated in Figure 1.
Figure 1. Baseline Experiment Topology
The workloads for single-stream, multi-stream, and connection migration are random plaintext generated by the Faker library (fak, 2024). For web page loading, we followed the methodology of Marx et al. (Marx et al., 2019) and Sander et al. (Sander et al., 2022), selecting 18 popular websites with varying numbers of resources and sizes shown in Figure 2. The websites were downloaded using Wget (wge, 2024) and hosted locally without external links on the Nginx server to eliminate network variability.
Figure 2. Website Sizes and Resource Counts
In the connection migration experiment, We extended Cloudflare’s Rust-based quiche client to support connection migration. Specifically, we took inspiration from the design of Andriod’s Connection Manager (Con, 2024) to detect changed interfaces/IP addresses, and initiate the connection migration process as described in (Iyengar and Thomson, 2022). For H2, we employed wget with and without Range Requests (ran, 2024), along with curl tests without Range Requests. Range requests in H2 facilitate fetching a range of data. This allows us to explore H2’s ability to pause and resume downloading in the change in connection scenario. The experiment simulates IP address changes after downloading over 2MB of a 5MB test file to mimic real-world network transitions, evaluating how each protocol maintains connection stability and download continuity during these transitions.
3.2. Proxy Experiment Setup
This experiment assesses the impact of introducing an Envoy reverse proxy between the server and client, with loss and delay applied between the proxy and the client. The reverse proxy breaks the connection between the client and the server and establishes a new connection with the server. We use BBR as the CCA for both QUIC and TCP since it is the default CCA for QUIC on Envoy. We disabled the proxy cache and retry features. The connection bandwidth is uniformly set at 10 Mbps across all links. We chose Envoy for its robust, production-grade capabilities and widespread adoption by major companies like Google, Microsoft, and Amazon. The network topology positions the Envoy between the client and server, as detailed in the accompanying Figure 3.
Figure 3. Proxy Experiment Topology
(a) 0% Loss, 0ms Delay
(b) 0% Loss, 50ms Delay
(c) 4% Loss, 50ms Delay
Figure 4. Single Stream Results. The times represent the mean values over 10 runs.
4. Experimental Evaluation
In this section, we present the results from our experiments and our analysis of the results.
4.1. Single Stream Experiment Evaluation
Figure 4 presents the results of our single stream experiments across various network conditions, highlighting the performance differences between H2 and H3.
Ideal Network Conditions Under ideal network conditions (10 Mbps bandwidth, 0% packet loss, 0 ms delay), both H2 and H3 exhibited similar performance, taking approximately 9 seconds to download a 10 MB file (see Figure 4(a)). The introduction of a proxy resulted in a minor increase in download time for both protocols, adding roughly 0-1 seconds. This proxy overhead was slightly more pronounced for H3 than for H2, likely due to Envoy Proxy’s experimental support for H3, compared to its mature, optimized support for H2.
Moderate Network Impairments In scenarios involving moderate impairments (10 Mbps bandwidth, 0% packet loss, and 50 ms delay), both H2 and H3 demonstrated comparable performance without the use of a proxy (Figure 4(b)). However, The introduction of a proxy caused slight variations in performance for H3 clients, such as curl and quiche, but overall, H3 maintained consistent performance with minimal impact from the proxy. In contrast, H2’s performance remained stable across different clients, suggesting that H2 implementations are more mature and better optimized for proxy environments. This indicates that further optimization is needed for some of the H3 clients to handle proxy-related challenges effectively.
In scenarios with loss only (10 Mbps bandwidth, 2-4% packet loss, 0 ms delay), both H2 and H3 exhibited comparable performance, with minimal impact from proxy use. This suggests that both protocols efficiently handle packet loss under these specific conditions.
Severe Network Impairments Under severe network conditions (10 Mbps bandwidth, 4% packet loss, and 50 ms delay), the performance differences between H2 and H3 became more pronounced (Figure 4(c)). H3 clients were able to complete the 10 MB file download in about 15 seconds, whereas H2 clients took approximately 106 seconds without a proxy. The introduction of a proxy using the BBR CCA significantly improved H2’s performance, reducing download times to around 10 seconds, representing a 90% improvement. Notably, other CCAs like Cubic and New Reno did not show this enhancement for H2, underlining the importance of CCA selection in impaired networks. However, it’s crucial to note that H2’s performance remains highly dependent on the chosen CCA, unlike H3, which is less impacted by such variations. H3’s performance remained consistent regardless of the CCA used, indicating that QUIC’s internal congestion control mechanisms are robust under such conditions.
Connection Establishment The QUIC connection starts with a handshake that merges the TCP 3-way handshake and the TLS 1.3 handshake into a single three-packet exchange. This process eliminates a full Round-Trip Time (RTT) in the QUIC startup phase, offering significant performance improvements for short sessions. While both protocols show similar connection establishment times under ideal network conditions, H3’s advantages become evident under impaired conditions (Figure 4(c)). The quicker connection establishment does not significantly impact performance in optimal conditions, but it becomes crucial in severely impaired networks.
Summary of Findings These experiments demonstrate that network conditions significantly influence the performance of both H2, with and without a proxy. H3’s superior performance in high-loss and high-latency environments is attributed to QUIC’s advanced congestion control and loss recovery mechanisms, which ensure accurate RTT measurements, efficient packet loss handling, and adaptive probe timeouts. In contrast, H2’s performance varies more significantly with network conditions and is heavily influenced by the chosen CCA.
Impact of CCAs In proxy scenarios, the use of BBR has shown to maintain high throughput and low latency by dynamically adjusting the congestion window based on real-time bandwidth and RTT estimations. This proactive congestion management is more effective than loss-based algorithms like Cubic, which only reduce the congestion window upon detecting packet loss. While BBR improves H2’s performance significantly in proxy-enhanced environments, this improvement is highly dependent on the correct CCA choice. In contrast, H3’s performance remains robust across different network conditions and CCAs, highlighting its superior consistency without relying on external configurations like BBR.
Proxy Influence and Implementation Maturity The slight overhead observed with proxies on H3 likely stems from the experimental support for H3 in Envoy Proxy, while H2 benefits from a more mature and optimized implementation. As H3 support within proxies matures, this overhead may decrease, potentially yielding further performance improvements. For now, H3’s performance remains robust across a variety of conditions, with the proxy introducing a more neutral impact compared to H2, which sees positive effects in proxy environments, especially when using BBR.
Figure 5. Multi Stream Results. Each point represents the highest Total Request Time among 20 test files for each run. A total of 10 runs were conducted for each loss level.
Stable IPIP ChangeStable IP With ProxyIP Change With Proxy
(a) 0% Loss, 0ms Delay
(b) 0% Loss, 50ms Delay
(c) 4% Loss, 50ms Delay
Figure 6. Connection Migration Results. Mean Times of 10 Runs with Standard Deviation Error Bars.
4.2. Multi-Stream Experiment Evaluation
The resulting graph in Figure 5 shows the performance of H2 and H3 under various loss levels with 20 streams. The data indicates that when the packet loss is below 8%, H2 and H3 exhibit comparable performance. However, when the packet loss reaches or exceeds 8%, H3 significantly outperforms H2, with the performance gap widening as the loss percentage increases. This demonstrates H3’s superior resilience to high packet loss conditions.
When downloading 20 1MB files concurrently, H3 consistently maintains stable download times, whereas H2’s download times vary significantly, particularly at a 12% loss level. This variability suggests that H2 struggles with higher loss levels, while H3’s performance remains consistent.
H3’s inherent mechanisms, such as QUIC’s ability to manage multiple streams within a single connection, contribute to its robustness. Each stream in QUIC independently manages its flow control and retransmission of lost data, making H3 less sensitive to increases in the number of streams. This approach effectively solves the HoL blocking problem in TCP. Overall, H3 is a more reliable choice compared to H2 in high-loss and multiplexing environments.
Analysis: The key advantage of H3 in multi-stream scenarios is rooted in its use of QUIC, which eliminates the HoL blocking problem commonly seen in TCP-based H2. In H2, packet loss in one stream can cause delays in other streams due to TCP’s sequential packet delivery requirement. This issue is exacerbated as the number of streams increases or under high packet loss conditions, leading to significant variability in download times. In contrast, QUIC’s stream independence allows each stream to handle its own packet loss and recovery, ensuring that packet loss in one stream does not impact others. This design enables H3 to maintain stable performance across multiple streams, even as packet loss increases.
Additionally, QUIC’s more sophisticated congestion control mechanisms, which include accurate RTT measurements and dynamic loss recovery, further enhance H3’s resilience. These mechanisms allow H3 to adapt quickly to changing network conditions, reducing the impact of packet loss and preventing the kind of performance degradation observed with H2.
Proxy Impact: The proxy has a neutral impact on H3, meaning it does not fundamentally alter H3’s performance. While H3 may see a minor impact under specific network conditions, these enhancements are minimal compared to H2, which sees a significant performance boost, particularly when using BBR. This makes H3 more predictable and stable in complex network environments, even with or without the use of a proxy.
4.3. Connection Migration Experiment Evaluation
Figure 6 displays the results of the connection migration experiments, comparing the performance of H3 and H2 clients under various network conditions. These results highlight the advantages of QUIC’s connection migration feature. The H3 client (quiche_h3) consistently exhibited lower total request times across all scenarios involving IP address changes.
Across all tested network conditions, H2 clients lacking the TCP range request capability (e.g., wget_h2_no_resume and curl_h2_no_resume) showed significantly higher total request times and were more adversely affected by IP changes and poor network conditions. However, when the TCP range request feature was employed (wget_h2_with_resume), H2’s performance became comparable to H3 in scenarios involving either packet loss or delay alone. While the TCP range request feature can be advantageous for specific use cases like file downloads, H3 is generally more suited for a broader range of applications, including real-time environments and the noisy network conditions often encountered in mobile scenarios.
Under conditions characterized by high packet loss and high delay, there were occasional instances where only 1 or 2 QUIC connection migrations failed. The results also indicate that H2, when used with a proxy, significantly outperforms direct H2, primarily due to the proxy’s use of the BBR CCA, as discussed in Section 4.1.
Analysis The primary advantage of H3’s connection migration feature is rooted in QUIC’s ability to seamlessly manage changes in network paths, such as those that occur when a device transitions between Wi-Fi and cellular networks. QUIC’s connection migration allows an ongoing session to continue without interruption, preserving the state of the connection even when the client’s IP address changes. This capability extends to proxy-enhanced environments, where H3 continues to maintain stable performance despite the additional complexity introduced by the proxy. In contrast, TCP-based H2 requires a new connection to be established following an IP change, leading to higher latency and increased total request times. This issue is especially noticeable in clients that do not support the TCP range request feature.
The results indicate that H3’s connection migration not only maintains the session but also minimizes total request times, which is essential for real-time applications where reducing delays is critical. This efficiency is a direct result of QUIC’s design, which includes mechanisms to quickly re-establish connection parameters and resume data transmission following an IP change. The fact that H3 rarely experiences failed migrations, even under high-loss and high-delay conditions, further underscores its robustness.
Impact of TCP Range Requests The use of TCP range requests in H2 mitigates some disadvantages associated with the absence of inherent connection migration. By enabling the resumption of downloads from the point of interruption, range requests provide a workaround to maintain connection continuity. However, this approach is more suited to static content downloads, where data can be easily segmented, rather than real-time data streams that require continuous and immediate transmission.
Role of Proxies and Congestion Control The significant performance improvement of H2 when using a proxy can be largely attributed to the implementation of the BBR CCA. BBR enhances throughput by dynamically adjusting to available bandwidth and reducing latency, which aids H2 in recovering more efficiently from connection interruptions. This illustrates the potential role of proxies in enhancing the performance of TCP-based protocols, albeit with added complexity. Conversely, H3’s inherent features reduce its reliance on proxies for performance optimization, making it a more suitable choice for environments where simplicity and robustness are paramount.
By leveraging QUIC’s native capabilities, H3 maintains a consistent performance advantage in dynamic and unreliable network conditions, making it a more reliable and effective option for scenarios that require robust connection migration.
H3 DirectH2 DirectH3 With ProxyH2 With Proxy
(a) 0% Loss, 0ms Delay
(b) 0% Loss, 50ms Delay
(c) 4% Loss, 50ms Delay
Figure 7. Web Page Loading Results. Mean Times of 10 Runs with Standard Deviation Error Bars.
$4.$ Web Page Loading Experiment Evaluation
In this experiment, we evaluated the performance of H2 and H3 protocols using the Speed Index as the primary metric for assessing user experience during web page loading. The Speed Index measures how quickly content appears on the user’s screen, with scores of 0-3.4 seconds indicating fast performance, 3.4-5.8 seconds indicating moderate performance, and scores over 5.8 seconds indicating slow performance.
Under baseline conditions (0% loss, and 0ms delay), as shown in Figure 7(a), the performance of H2 and H3 varied across different websites. Overall, H2 and H3 were comparable, with both protocols predominantly achieving a Speed Index within the ”fast” category. When a proxy was introduced, the impact on performance varied: some websites experienced better performance while others showed slight degradation. Notably, more than half of web pages performed better with a proxy under no loss and no delay conditions for both H2 and H3. When loss and delay were introduced, nearly all web pages achieved better or the same performance compared to direct requests.
In the impaired network scenario (4% packet loss, and 50ms delay), as shown in Figure 7(c), H2’s performance significantly deteriorated, with many websites falling into the ”moderate” or ”slow” categories. In contrast, H3 demonstrated greater resilience, maintaining better performance with most websites remaining in the ”fast” or lower ”moderate” categories. While H2’s performance improved with the use of a proxy and BBR, this improvement is largely due to the dependency on BBR, as discussed in §4.1. H3, by comparison, exhibits strong performance in handling network impairments without needing reliance on external elements like specific CCAs, further demonstrating its robustness. Additionally, larger websites with more resources generally had higher Speed Index scores, particularly under impaired conditions. In the less impaired network scenario (2% packet loss, and 25ms delay), we observed similar conclusions to the 4% loss and 50ms delay scenario, but with lower Speed Index scores.
Analysis: The observed differences in web page loading performance between H2 and H3 can be attributed to QUIC’s inherent advantages in handling multiplexed connections. In H2, multiplexing over a single TCP connection can still suffer from HoL blocking if packet loss occurs, as all streams within the connection are affected by the loss. This is particularly problematic for loading web pages with many resources, where the delay in one stream can delay others. In contrast, QUIC, used by H3, avoids HoL blocking by managing each stream independently, allowing lost packets to be retransmitted without affecting the delivery of other streams. This results in a smoother loading experience and a better Speed Index, even under challenging network conditions.
Efficiency in Multiplexing: H3’s better handling of multiplexing not only reduces the impact of packet loss but also optimizes the utilization of available bandwidth. By maintaining separate flow control and loss recovery for each stream, QUIC ensures that multiple resources can be fetched concurrently without waiting for others to complete. This capability is crucial for modern web pages, which often consist of numerous small objects that need to be loaded simultaneously. The improved Speed Index for H3 in impaired conditions demonstrates that QUIC’s design is more suited to real-world web traffic patterns, which frequently involve high concurrency.
Proxy Impact and Congestion Control: The use of a proxy, particularly with BBR as the CCA, helps mitigate the effects of network impairments by optimizing throughput and reducing latency. The proxy’s positive impact on H2 performance highlights how BBR’s proactive congestion management can alleviate some of the drawbacks of TCP’s loss-based congestion control, reducing the frequency and duration of HoL blocking events. For H3, the more consistent performance with or without a proxy underscores QUIC’s capability to manage congestion and packet loss internally, making it less reliant on external optimizations provided by proxies.
By leveraging its advanced multiplexing and independent stream management capabilities, H3 delivers a more resilient and efficient web browsing experience, making it particularly well-suited for environments with variable and challenging network conditions.
4.5. General Observations
Across all experiments, the results consistently highlight the superior performance of H3 over H2 in a variety of network conditions, particularly in scenarios involving packet loss, delay, and high concurrency. The fundamental architectural differences between H2 and H3, particularly the underlying transport protocols (TCP for H2 and QUIC for H3), play a crucial role in these performance disparities.
1. QUIC’s Robustness to Network Impairments: One of the most significant advantages of H3 is its resilience in impaired network environments. QUIC’s design, which includes features like connection migration, independent stream management, and advanced loss recovery mechanisms, ensures that H3 can maintain stable performance even under high packet loss and latency conditions.
2. Efficient Multiplexing and Avoidance of HoL Blocking: A recurring theme in the experiments is H3’s ability to handle multiple streams efficiently. Unlike H2, where packet loss in one stream can cause delays in others due to TCP’s sequential nature, QUIC enables each stream to manage its flow control and retransmission independently. This feature minimizes the impact of packet loss on overall performance, making H3 particularly well-suited for web page loading scenarios with numerous small resources. The ability to avoid HoL blocking directly contributes to H3’s better Speed Index scores and consistent performance in multi-stream experiments.
3. Role of Congestion Control: The experiments also underscore the impact of CCAs on protocol performance. While both H2 and H3 can benefit from advanced CCAs like BBR, H3’s internal congestion control mechanisms appear more adept at handling varying network conditions. The use of BBR in proxies significantly boosts H2’s performance by reducing latency and optimizing throughput. However, H3’s inherent design makes it less reliant on external congestion management.
4. Connection Migration Capability: QUIC’s connection migration feature offers a clear advantage in scenarios involving network changes, such as mobile environments where devices frequently switch between Wi-Fi and cellular networks. This capability allows QUIC to seamlessly maintain connection continuity and low request times despite IP address changes. Notably, QUIC’s connection migration remains robust even in proxy-enhanced environments, further demonstrating its resilience and suitability for dynamic network architectures. In contrast, TCP lacks this flexibility, resulting in higher latency and connection interruptions, even when augmented with features like TCP range requests, which are only beneficial for static downloads rather than real-time applications.
5. Proxy Usage and Performance Impact: The experiments reveal that proxies can significantly enhance H2 performance, especially in impaired networks, by leveraging advanced CCAs and optimizing connection handling. However, the impact on H3 is less pronounced, suggesting that QUIC’s internal optimizations are sufficient to manage most network conditions effectively. This makes H3 a more versatile choice in various deployment scenarios, reducing the dependency on complex proxy configurations.
6. Practical Implications: The findings suggest that H3 is better equipped to handle modern web traffic patterns, characterized by high concurrency, frequent network changes, and varying quality of service. H3’s design aligns well with the requirements of real-time applications, mobile browsing, and environments with challenging network conditions. As such, the adoption of H3 can lead to improved user experiences, reduced latency, and more efficient use of network resources.
Overall, the comprehensive evaluation across different scenarios demonstrates that QUIC’s advanced features provide H3 with a significant edge over H2 in terms of performance, reliability, and adaptability. These findings have broader implications for the future of web protocols, particularly as the demand for faster, more reliable internet communication continues to grow.
5. Future Work
Future research should focus on exploring the interaction of different CCAs with HTTP/3 across various network environments. Beyond BBR, testing algorithms like Cubic, NewReno, and newer ones could reveal how they affect QUIC’s performance, offering insights into optimizing configurations for specific use cases such as real-time applications, video streaming, and mobile networks. Examining these factors could improve the efficiency and adaptability of H3 in diverse scenarios.
Additionally, investigating the role of proxies in optimizing web protocols, especially in edge computing and CDN environments, would be valuable. Future work could enhance H3 support in proxies and develop specific optimizations for QUIC, particularly for high-latency or lossy networks. Moreover, studies on security and privacy implications related to QUIC’s fast connection establishment and migration could assess their impact on data integrity and confidentiality in scenarios like man-in-the-middle attacks or network handoffs. This research is essential for securing H3 in sensitive environments and large-scale deployments under real-world traffic conditions.
6. Conclusion
This study demonstrates H3’s superior performance over H2, especially in challenging network conditions with high packet loss and latency. H3’s features, such as connection migration and multiplexing, allow it to maintain stable performance, even with a proxy. Unlike H2, which relies on the right CCA, H3 delivers consistent reliability across diverse scenarios.
The main takeaway is H3’s robustness in both proxy and non-proxy environments, making it a more dependable choice for modern web communication. H3 consistently outperforms H2 without the need for external optimizations.
In conclusion, H3’s design ensures resilient and efficient operation in both stable and unstable networks, solidifying its role as the preferred protocol for future web communication.
Appendix A Research Methods
A.1. Clients and Server Protocols and Implementations
The clients and server protocols and implementations show in Table 2.
Table 2. Clients and Server Protocols and Implementations
Client Protocol Language Version
Proxygen (pro, 2024) QUIC/H3 C++ v2024.03.25.00
Curl (cur, 2024) QUIC/H3 C OpenSSL (Ope, 2024): v3.3.1, nghttp3 (ngh, 2024): v1.1.0, Curl:8.7.2-DEV
Curl (cur, 2024) TCP/H2 C nghttp2/1.40.0 (nghttp2 Team, 2024)
Chrome (chr, 2024) TCP/H2 C, C++ Node.js (nod, 2024): v18.17.1, Puppeteer (Pup, 2024): 22.6.1, HeadlessChrome /123.0.6312.58, Puppeteer-har: 1.1.2
Chrome (chr, 2024) QUIC/H3 C, C++ Node.js: v18.17.1, Puppeteer: 22.6.1, HeadlessChrome /123.0.6312.58, Puppeteer-har: 1.1.2
Wget (wge, 2024) TCP/H2 C Wget:1.20.3-1ubuntu2
Quiche QUIC/H3 Rust Quiche: 0.20.0-47-g098664eb
aioquic (aio, 2024) QUIC/H3 Python Aioquic:1.0.0-2-ga6d91fb
Nginx (ngi, 2024) TCP/H2 QUIC/H3 C nginx/1.25.5